OpenStudy (anonymous):

An 18 foot ladder is leaned against a wall. If the base of the ladder is 7 feet from the wall, how high up on the wall does the ladder reach?...Please show work.. thanks ;)

OpenStudy (anonymous):

18 is hypotenuse, 7 is base. find height a^2+b^2=c^2 7^2+b^2=18^2 49+b^2=324 b^2=275 b=16.583123951777 Height =b. Or 16.583123951777